steam at 7bar and 250 C enters a pipeline and flow along it at constant pressure. if the stream rejects heat steadily to the surroundings. at what temperature will droplets of water begin to form in the vapour ?. using steady flow energy equation and neglecting changes in the velocity. calculate the heat rejected per kilogram of steam flow
1
Expert's answer
2020-04-02T09:32:04-0400

Here, pressure of steam is given as 7 bar and it is flowing in pipe at temp of 250oC

And the steam is ejected to surrounding at constant pressure and forms water droplet.

The temperature at which it changes to water at constant pressure is defined as saturation temperature.

From the steam table for constant pressure line we can take value ,



We get the saturation temperature or the temperature at which water droplets forming is

Tsat= 1650C


Now , for finding the heat rejected per kg we can use stead flow energy equation



h1 enthalpy at 150 oC=2762 kJ/kg and h2 ,enthalpy at 250 oC=2954


Here heat rejected= h2-h1=2954-2762= 191 kJ/kg